OpenStudy (anonymous):

find the deivative of f(x)=1/x^2

OpenStudy (anonymous):

it is -2/x^3

OpenStudy (anonymous):

Note that 1/x^2 = x^(-2).

OpenStudy (anonymous):

derivative*

OpenStudy (anonymous):

this means \[f'(x)=-\frac{2}{x^3}\]

OpenStudy (anonymous):

Hey satellite how can I enter LaTeX in here?

OpenStudy (anonymous):

how did you get there?

OpenStudy (anonymous):

Using the rule for differentiating polynomials.

OpenStudy (anonymous):

d/dx ( x^n ) = n*x^(n-1)

OpenStudy (anonymous):

This holds for all real n.

OpenStudy (anonymous):

i have to use f(x + dx) - f(x)/dx

OpenStudy (anonymous):

Don't forget your parenthesis in the numerator!

OpenStudy (anonymous):

yes all of it is in the numerator

OpenStudy (anonymous):

you enter latex via "\[" to open and "\]" to close

OpenStudy (anonymous):

Then you need to use the binomial theorem to prove the identity for general polynomials.

OpenStudy (anonymous):

dang didn't work @Ymanifold come to chat and i can show you there

OpenStudy (anonymous):

I get the idea satellite

OpenStudy (anonymous):

we haven't learn that

OpenStudy (anonymous):

\[ f(x) \]

OpenStudy (anonymous):

there you go and if you right click you can see the source

OpenStudy (anonymous):

ahhhhhh frustration!

OpenStudy (anonymous):

\[\int_0^{\infty}e^{-x^2}dx\]

OpenStudy (anonymous):

thanks satellite

OpenStudy (anonymous):

welcome @amber do you have to find this derivative from scratch? using the definition?

OpenStudy (anonymous):

yea thats what she has to do

OpenStudy (anonymous):

be good practice for your latex. you gonna write it out?

OpenStudy (anonymous):

okay i was with you until the end

OpenStudy (anonymous):

That is the geometric sum

OpenStudy (anonymous):

You need it to expand 1/x

OpenStudy (anonymous):

ahhh just forget it ill ask my teacher but thanks for trying!

OpenStudy (anonymous):

all i am doing is finding the slope of the tangent curve

OpenStudy (anonymous):

What I'm saying is: Use \[ \frac{1}{1-x}=\sum^\infty_{k=0} x^k \]

OpenStudy (anonymous):

\[\frac{f(x+h)-f(x)}{h}=\frac{\frac{1}{(x+h)^2}-\frac{1}{x^2}}{h}\] \[=\frac{x^2-(x+h)^2}{x^2(x+h)^2h}\] \[=\frac{x^2-(x^2+2xh+h^2)}{x^2(x+h)^2h}\] \[=\frac{-2xh+h^2}{x^2(x+h)^2h}\] \[=\frac{h(-2x+h)}{x^2(x+h)^2h}\] \[=\frac{-2x+h}{x^2(x+h)^2}\] now you no longer have and h as a factor in the denominator, so take the limit by replacing h by zero and get the answer

OpenStudy (anonymous):

Thats much easier :D.

OpenStudy (anonymous):

well it is just a bunch of ugly algebra, the gimmick of which is to get the h to cancel top and bottom so you can take the limit

OpenStudy (anonymous):

ok i kinda got that! but i had (-2xh - h^2)/x^2(x+h)^2

OpenStudy (anonymous):

NEVERMIND!!! i had that the whole time just was mixed up THANKS! ((:

OpenStudy (anonymous):

lol!

OpenStudy (anonymous):

maybe you forgot the h in the denominator, or didn't factor and cancel. good job though

OpenStudy (anonymous):

yea i did forget the h